# Rethinking the Integration of Prediction and Planning in Deep Learning-Based Automated Driving Systems: A Review

### v2

@@ 3.1 Scene Representation
An alternative way to achieve viewpoint invariance are pairwise relative coordinate systems introduced in GoRela [97], which describe the relation between agents instead of using a fixed coordinate frame. This allows the computation of the relation between static objects offline and greatly reduces computation. PEP [94] further reduces computation by using center-relative instead of pairwise relative coordinates, i.e. shifting the coordinate frame into the mean position of all processed agents before each transformation.
Another choice to increase robustness are Frenet representations that decompose positions into progress along and lateral offset from a lane [98, 99]. Either a single Frenet frame defined by the target vehicle is used for all agents [100] or each agent is represented in an individual coordinate frame [84, 101].
2016201720182019202020212022202320241988DESIRE [8]Fast and Furious $\uparrow$ [102]R2P2 $\downarrow$ [103]MTP [54]IntentNet [49]MATF [89]MultiPath $\uparrow$ [56]MFP [88]CoverNet $\downarrow$ [55]Trajectron++ [40]VectorNet [74]LaneGCN $\uparrow$ [41]TNT $\downarrow$ [75]mmTransformer [76]HOME $\uparrow$ [104]AgentFormer $\downarrow$ [105]DenseTNT $\uparrow$ [106]GOHOME $\downarrow$ [84]MultiPath++ [38]SceneTransformer $\uparrow$ [6]HEAT $\downarrow$ [107]HiVT [108]Wayformer [109]Motion Transformer [110]GoRela [97]LaFormer [111]MotionDiffuser [112]Trajeglish $\uparrow$ [113]BAT $\downarrow$ [114]Scene Representation (3.1)RasterizedSparseMixedInteraction Model (3.2)RNNCNNGNN / AttentionTransformerTrajectory Decoding (3.3)Trajectory RegressionAnchor Trajectory ClassificationEndpoint and CompletionInitial Trajectory and RefineAuto-Regressive FormulationPrediction2016201720182019202020212022202320241988ALVINN [9]Deepdriving [115]End-to-End-CNN [19]FCN-LSTM [116]CIL [117]Camera360 [118]Seg2WP [22]CAL [119]DeepImitative [62]DOCP [10]ChauffeurNet [58]DILP [63]CILRS [64]NMP [61]UD-CIL $\uparrow$ [23]NEAT $\uparrow$ [120]JLTP $\downarrow$ [121]LBC [67]LSD [122]PILOT [123]Transfuser [20]UrbanDriver $\downarrow$ [24]LAV [124]PlanT [125]PDM-Open [126]Input Space (4.1)LatentInterpretableOutput Space (4.2)ActionTrajectoryGoal Conditioning (4.3)NoneSubmodulesInput FeatureRoute CostRoute AttentionPlanning Paradigm (4.4)RegressionCost FunctionHybridPlanning2016201720182019202020212022202320241988PAC [127]HRI [128]PRECOG [65]PiP [129]DSDNet $\uparrow$ [68]P3 $\downarrow$ [66]LookOut $\uparrow$ [69]MP3 $\downarrow$ [70]SafetyNet [130]DIPP [7]SafePathNet [131]GameFormer $\uparrow$ [132]PDM-Hybrid $\uparrow$ [126]UniAD $\downarrow$ [133]TPP [134]VAD $\downarrow$ [135]DTPP [136]Ego-Agents Relation (5.2)UnknownRobot LeaderHuman LeaderJoint PlanningCo-LeaderSafety & Contingency (5.3)No CostMarginalizedWorst-CaseJoint CostContingentGoal Conditioning (4.3)NoneSubmodulesInput FeatureRoute CostRoute AttentionPlanning Paradigm (4.4)RegressionCost FunctionHybridManual IPPS
Figure 3: Contributions that got attention in the research fields of DL-based prediction, planning, and manual IPPS for automated driving. Nodes, stem, and label of each entry encode different aspects that are explained below the timelines. Works marked with $\uparrow$ and $\downarrow$ were slightly moved up and down the timeline, respectively, to avoid overlaps.
3.2 Interaction Modeling
A decisive aspect of prediction is to model the interaction between traffic agents represented in the scene [4, 137]. This includes interaction with map elements and other traffic agents. Modeling the relation of vehicles to static map elements like lane markings, traffic signs, etc. is important to understand feasible corridors in which trajectories could be localized. Interaction with other traffic agents are crucial to understanding social interaction.
RNNs. A few early prediction models like DESIRE [8] or Trajectron++ [40] combine RNNs with an aggregation operator like spatial pooling or attention. They either extract temporal per-agent features and then aggregate them across all agents or aggregate information first and extract joint temporal features afterwards.
CNNs. Alternatively, Fast and Furious [102], MTP [54], and Multipath [56] applied 2D convolutions to implicitly capture interaction within the kernel size. Compared to sequence processing-based approaches, higher importance is assigned to spatial interaction.
GNNs and Graph-Attention explicitly model interactions between individual agents by combining multiple agents’ features [138] and processing them with graph convolution operators [139, 73, 140, 141] or graph attention [142, 143, 144, 145] to aggregate information. Prediction networks like VectorNet [74], LaneGCN [41], TNT [75], HOME [104], DenseTNT [106], GOHOME [84], Multipath++ [38], LaFormer [111], and BAT [114] are examples of this development.
Transformers [83] have been widely adopted in prediction: InteractionTransformer [146], mmTransformer [76], AgentFormer [105], SceneTransformer [6], HiVT [108], Wayformer [109], Motion Transformer [110], MotionDiffuser [112], Trajeglish [113] and many further works [147, 148, 149, 150, 110, 151, 152, 153, 154, 155, 156] exploit the global receptive field and attention mechanism of Transformers. All SVs can be predicted simultaneously [76], and the impact of vehicles on each others’ behaviors across different timesteps can be modeled with attention layers [105]. Attention layers can be applied along the spatial and temporal dimension jointly [105], or individually [157]. The latter can be divided into sequential [158, 159] and interleaved [6, 109] architectures. Experiments with Wayformer [109] and SceneTransformer [6] show that joint and interleaved processing outperform sequential approaches. As, joint processing causes higher computational burden, the interleaved paradigm balances best between performance and computational demand.
The timeline in Fig. 3 reflects how interaction models changed together with the scene representation. After CNNs and RNNs were applied to rasterized scene representations in early works, GNNs, attention, and Transformers took over since the advent of sparse representations.
Across all architectures, the prevalent concept is to extend interaction from a local to a global scale. HiVT’s ablation study [108] shows that both, local and global attention, are beneficial for prediction and that hierarchical attention yields the best performance.
3.3 Trajectory Decoding
The final step of prediction is to generate trajectories from latent features. As introduced in Sec. 2, this step models a probability distribution $P_{\textrm{pred}}=P(\overline{\textbf{Y}}|\overline{\textbf{X}},I)$ over possible future trajectories $\overline{\textbf{Y}}$ given observation $\overline{\textbf{X}}$ and scene context $I$. Here, we focus on two major aspects of trajectory decoding: The modeling of multimodality and the decoding principle.
Multimodality. The intention of SVs is inherently uncertain, making $P_{\textrm{pred}}$ multimodal [160]. Therefore, $P_{\textrm{pred}}$ is commonly either expressed as an explicit density function or estimated with a discrete trajectory set.
@@ 3.3 Trajectory Decoding
Explicit density functions. An intuitive approach to prediction is to explicitly model the probability density function of $P_{\textrm{pred}}$ [61, 66, 69, 70]. Continuous distributions like bi-variate Gaussian distributions [161], Gaussian Mixture Models [110, 162, 132], or discrete rasterized heatmaps [104, 84] can be found in many prediction works. Alternatively, object-agnostic representations like occupancy maps [53, 163], or flow fields [164, 165, 70] can be used. These representations handle multimodality implicitly [61] and were shown to be more robust against perturbations [62]. However, since no discrete trajectories are decoded, it is difficult to compare such representations with expert logs for performance evaluation.
@@
Discrete trajectory sets are either created by sampling from an intermediate distribution or by the model design. To sample a diverse trajectory set from an intermediate distribution, generative models such as Generative Adversarial Networks (GANs) [166], Conditional Variational Autoencoders (CVAEs) [167], normalizing flows [168] or denoising diffusion [112], can be used [8, 89, 169, 170, 171]. Alternatively, recent methods apply several prediction heads to the features extracted by a common backbone to decode a set of diverse trajectories [38, 41, 172, 109]. Other ways to obtain diverse predictions are training loss functions [75], entropy maximization [39, 173], variance-based non-maximum suppression [174], greedy goal sampling [55, 97], divide and conquer strategy [175], evenly spaced goal states [176], or with pre-defined anchor trajectories [56, 177]. Nevertheless, a guarantee of coverage can only be given by specifying high-level behaviors [124] or regions of the map that must be covered by at least one prediction [76].
Decoding principles. We classify the decoding principles into one-shot, two-shot, and n-shot methods as depicted in Fig. 4. While it is a common practice in planning to decode trajectories as a sequence of actions that translate into Cartesian space using a kinematic model (cf. Sec. 4.2), only a few works do so in prediction [178, 179]. Instead, future waypoints are predicted directly from latent features [41, 38, 111, 109, 180].
Single-shot approaches are split into trajectory regression and anchor trajectory classification.
Trajectory regression directly decodes the latent features to a trajectory using a NN, typically an MLP [54, 74, 41, 38, 6, 107, 108, 109]. Especially Transformer-based architectures use direct trajectory regression decoding (cf. Fig. 3). However, these methods are prone to predict trajectories that leave the road, cut corners, or are kinematically infeasible [97]. Specifically designed loss functions for offroad driving [181, 182, 39, 183, 184], driving direction compliance [185] or distance to the centerline [186] can alleviate the problem.
Anchor trajectory classification states prediction as a classification task. Thus, a probability is estimated for a set of pre-defined trajectories [55, 177]. Defining anchor trajectories can ensure the feasibility of the prediction and allows to impose hard kinematic constraints. However, rare trajectories might not be covered by the anchors.
Two-shot methods structure trajectory decoding into meaningful subtasks that increase explainability and are more flexible than anchor trajectory classification. They comprise endpoint and completion as well as initial trajectory and refine strategies.
Endpoint and completion methods are based on the assumption that a trajectory is mostly defined by its endpoint [25]. The endpoint can either be regressed [104][84] or classified from a pre-defined set [75][106][97]. Usually, classification-based methods additionally regress an offset of the classified endpoint [75]. HOME [104] estimates the endpoint probability distribution as a rasterized map, samples potential endpoints, and then regresses the trajectory using an MLP.
Initial trajectory and refine approaches combine the ideas of anchor trajectory classification and endpoint and completion. Instead of an endpoint, an initial trajectory is classified or generated. Then, a refinement step adapts the trajectory, for example, by offset regression for each predicted waypoint [102, 56, 111]. TPNet [187] and DCMS [188] first regress an endpoint, then generate trajectory proposals and finally refine them.
N-shot methods are either initial trajectory and refine or autoregressive formulations.
Initial trajectory and refine strategies in their n-shot variants are similar to the two-shot version. The only difference lies in the refinement step, which is realized as a recurrent optimization [8, 112]. For example, MotionDiffuser [112] samples a noisy trajectory and repeatedly applies denoising diffusion to refine the trajectory.
Autoregressive approaches use recurrent decoding to perform a stepwise roll-out into the future. Iteratively, the next waypoint is predicted, and information on the resulting scene is added to the latent features [103, 189, 40, 76, 105].
Autoregressive decoding can help to enforce high-quality social interactions and scene understanding [97] but might lead to compounding errors since predicted information is used to make further predictions [190]. Further, the computational effort and computation time of these methods are usually higher than for one-shot or two-shot methods.
Fig. 3 shows the development of trajectory decoding using the categories defined in this Section. While endpoint and complete decoding gained popularity, auto-regressive formulations were recently used less often, assumingly due to computation time deficiencies. Trajectory regression decoding was always used but became prevalent with the rise of Transformer-based interaction modeling in prediction.
Figure 4: Principles of trajectory decoding.
3.4 Benchmarks
Various prediction datasets provide training, validation, and test splits for benchmarking. The most prominent ones are nuScenes [191], WOMD [192], Argoverse [193], Argoverse2 [194], inD [195], highD [196], rounD [197], exiD [198], and Interaction [199]. Evaluation is carried out by comparing the predicted trajectory of an SV to its ground truth future trajectory. To account for the inherent multimodality benchmarks are commonly based on a winner-takes-all evaluation. This means the model predicts a fixed number of output trajectories, e.g., six [194] or ten [191], and the evaluation takes only the best one among them into account. Alternatively, predictors that output a parametric distribution are often evaluated based on the Negative log likelihood (NLL) of the ground truth under the predicted distribution. This assessment also factors in the uncertainty predicted by the model and evaluates the whole distribution instead of only a single trajectory.
4 Planning
The planning task is to find a trajectory $Y_{\textrm{EV}}$ for the EV that factors in safety, comfort, kinematic feasibility, and goal-directedness based on the observations $X_{\textrm{EV}}$ and $X_{\textrm{SV}}$ as well as additional context $I$ and optionally $\overline{\textbf{Y}}_{\textrm{SV}}$ (cf. Sec. 2). In this section, we will first give a comprehensive overview of common input (Sec. 4.1) and output representations (Sec. 4.2). Then, we address goal-conditioning in Sec. 4.3 before we categorize existing works and discuss common paradigms in Sec. 4.4. Finally, we briefly describe existing benchmarks in Sec. 4.5.
4.1 Input Representation
We classify planning input representations into interpretable intermediate representations and latent features.
Interpretable intermediate representations such as rasterized or sparse scene representations (that were discussed thoroughly in Sec. 3.1) are common inputs to planning components in modular ADSs. Thus, many works exist that build on rasterized [58, 67, 124, 200] or sparse inputs [125, 131, 201, 130, 24]. Similar to prediction, a clear trend can be observed from rasters to sparse inputs.
Latent features. E2E ADSs directly map sensor information to actions (cf. Fig. 1[9, 202, 20]. Thus, instead of $X_{\textrm{EV}}$, $\overline{\textbf{X}}_{\textrm{SV}}$, and $I$, the planner’s inputs are latent features.
@@ 4.1 Input Representation
The benefit of using learned latent intermediate representations is that no manual engineering effort is needed to design the interfaces between different modules. The main drawback of hand-engineered interfaces is that they have to apply to the long-tailed distribution of potential driving scenarios, and every piece of information discarded by the interface cannot be recovered. For instance, if a vehicle is currently occluded, but its presence can be inferred from the reflection of its headlight in another car, bounding box representations [58], occupancy fields [70], or affordances [115, 119] will not be able to propagate this information to downstream modules [203, 133]. In contrast, latent feature representations allow uncertainty propagation, enabling downstream modules to compensate for errors at earlier stages, such as undetected vehicles [61].
The main disadvantage of latent feature representations is the lack of interpretability. In case of failures, evaluating which part of the system caused the error is extremely difficult. This aggravates tuning and debugging of pure E2E systems. Interpretable E2E systems seek to compensate for this by producing additional intermediate representations [66]. In contrast to modular stacks, these interpretable intermediate outputs are not used by the planner. Instead, they are only used for additional supervision and model introspection.
The timeline in Fig. 3 reveals that interpretable input representations are becoming increasingly popular compared to latent input features.
4.2 Output representation
The planning output is typically represented as a sequence of future states or control actions.
Future states are given by a sequence of future 2D position and heading angles representing the planned trajectory of the EV over time [22, 62, 58]. This trajectory is then passed to a downstream controller tasked with tracking it [22]. The trajectory representation is well interpretable and can be easily checked for collisions, traffic law violations, or deviations from the driveable surface [129, 68]. Moreover, it is agnostic of the kinematic parameters of the controlled vehicle, enabling generalization to other vehicles. However, the ultimate driving performance also depends on the downstream controller and how well the planner and controller are tuned for each other. Validating and testing the system can require iterations between both components, which slows down development.
Future actions are an alternative representation of planned behavior. This is mostly adopted by E2E ADSs that regress acceleration and steering angle without the intermediate trajectory representation [117, 118, 119]. Besides ensuring kinematic feasibility, it can promote comfort as it is directly related to the magnitude of the actions.
However, the resulting behavior depends on the controlled vehicle’s individual dynamics [204, 19]. Hence, generalization to other vehicles is limited.
Fig. 3 depicts that early E2E ADSs planned in actions, while trajectory outputs recently gained popularity.
4.3 Goal Conditioning
The overall task of ADSs is to navigate safely toward a destination. Hence, goal-directedness is one main criterion that determines the suitability of a potential plan. Lane-level route information can be provided to the planner by a navigation system. Popular benchmarks provide route information through a high-level command such as ’turn left’ or ’go straight’ [205], through sparse goal points along the route [205] or by annotating lanes within the map as on-route or off-route [206]. For example, in the widely used Carla simulator [205], goal locations are sparsely sampled along the route, and in each planning step, the closest goal location is provided to the planner. We define four categories to describe how this goal information is incorporated into the planning algorithm: input features, separate submodules, routing cost, and route attention. While the first two were originally introduced in [117], the latter were proposed more recently [201, 207, 126, 208].
Input features are the most straightforward way to incorporate goal information and have been widely adopted [120, 20, 124, 209, 210, 24, 63, 64, 118, 23, 122]. Different methods exist to pass this input to the planning module depending on how the route information is represented. If the lanes within the map information are annotated as on-route or off-route, then this information can be used as a separate semantic channel in raster maps [58]. Similarly, vector inputs describing the lane centerlines can have an additional flag reflecting this information [201].
High-level commands were used as input feature in LBC [124], LAV [67], and AD-MLP [208]. It was found that inputting high-level intention repeatedly at different stages of the network increases its generalization capability [23]. Transfuser [20] leverages sparse goal locations as 2D features in the final trajectory decoding step.
While using goal information as an input feature is simple, goal compliance is not guaranteed. On the contrary, [201, 117] demonstrated that the model could ignore this additional input and instead rely on other potentially spurious correlations, such as between the current and the future kinematic state.
Separate submodules are used only with high-level commands, which serve as a switch between command-specific submodules [117, 70]. Balancing the dataset to prevent it from being dominated by a single driving mode, such as lane following, is avoided in this setup, as each block is only trained on a single high-level command. However, this approach requires a fixed number of high-level commands to be defined in advance.
Routing cost. A third option is to plan trajectories by optimizing a cost term for routing [66, 69, 62]. Such cost functions can measure the remaining distance to a goal location [62], the progress along or distance to an on-route lane [66]. Goal-directedness is balanced with further competing objectives such as safety and comfort. Thus, the planner can flexibly react to difficult situations, e.g., it could drive off the road to prevent a collision. However, tradeoffs between safety and comfort versus progress can be undesirable.
Route attention applies spatial attention to relevant parts of the map, i.e., the route. This can be achieved by removing the off-route portion of the map from the input features, e.g., PDM-Open [126] and PEP [94] discard almost the entire map information and only keep points along a lane centerline on the route. Similarly, GC-PGP [201] removes off-route nodes from the lane graph after information aggregation and TPP [134] and DTPP [136] discard all EV trajectory proposals that do not comply with the route.
Fig. 3 shows that early models did not condition their plan on a goal at all. Afterward, submodules and input features dominated. While using input features is easy to implement, it is less effective, whereas submodules represent a major architectural constraint. The most interpretable method is using a routing cost function. Recent models introduced route attention, which is more flexible. Both route attention and routing cost allow to condition on goals that are not reflected in the training dataset, thus improving generalization.
4.4 Planning Paradigms
@@
Recall that in Eqn. 5 we defined planning to be a function $f$ mapping from observational inputs $X_{\textrm{EV}}$, $\overline{\textbf{X}}_{\textrm{SV}}$, $I$ and predictions $\overline{\textbf{Y}}_{\textrm{SV}}$ to a trajectory $Y_{\textrm{EV}}$. In the following, we focus on the planning function $f$. To this end, we decompose $f$ into two parts: A proposal generator $g$, yielding a set of multiple potentially suitable trajectories $\hat{Y}^{(i)}_{\textrm{EV}},i=1,..,N_{\textrm{proposals}}$ and a proposal selector $h$ that selects the final plan $Y_{\textrm{EV}}$ among the proposed set. Hence, in this work, we propose to represent the planning function $f$ as
@@
Cost function optimization. As planning aims to find a trajectory that optimizes objectives such as safety, comfort, and progress, it is a natural approach to design and optimize a cost function that balances these potentially conflicting objectives. Cost function-based planning methods [121, 61, 211] rely entirely on the selection function $h$ to find a suitable trajectory, i.e.,
@@
where $c$ is a cost function of the planned trajectory. Consequently, the proposal generator $g$ may randomly sample feasible motion profiles [61] or obtain them by clustering real-world expert demonstrations [70]. Traditional non-learning-based methods rely on hand-crafted cost functions [212, 213, 214, 215, 216, 217, 218, 123]. However, designing a cost function that effectively balances these goals and generalizes to the long-tailed distribution of possible driving scenarios is challenging. Learning-based methods aim to address this by learning a cost function directly from expert demonstrations [219, 220, 103]. Similarly, [119, 115] learn to regress driving-relevant affordances such as the distance to the leading vehicle or the heading angle relative to the lane, based on which a downstream controller infers driving commands.
The learned cost function can be non-parametric [61, 221, 222] or comprise hand-designed sub-costs of which only the weights are learned [69, 70, 66, 121]. The non-parametric cost function of the NMP [61] assigns a cost to each potential future EV position within the planning horizon and is represented as a 3D-BEV-Tensor. In contrast, hand-designed sub-costs commonly evaluate safety, comfort, traffic rule compliance, and progress [121, 70, 66].
Regression approaches rely entirely on the proposal generator $g$. The selection function $h$ is the identity, and all burden is placed on the proposal generator $g$, which only generates a single proposal $Y_{\textrm{EV}}=\hat{Y}^{(1)}_{\textrm{EV}}$. E2E methods fall into this category. Regression approaches can learn to directly predict future waypoints [201, 20, 24] or actions [9] by supervised learning [223]. This form of imitation learning, i.e., supervised learning from expert data, is called behavior cloning (BC) [224, 190, 225]. This approach makes the manual design of the cost function obsolete. However, BC models are vulnerable to distributional shifts, which occur when the model reaches states not covered by the training data. Besides, regressed trajectories are generally not guaranteed to be kinematically feasible. UrbanDriver [24] counters this by training the model directly in a closed-loop simulation. However, this requires the entire simulator to be differentiable. CW-ERM [226] rolls out the policy in a non-differentiable simulation to identify important training scenarios and assign sample weights.

### v3

@@ III-A3 Coordinate Frame
An alternative way to achieve viewpoint invariance is the pairwise relative coordinate system introduced in GoRela [94], which describes the relation between agents instead of using a fixed coordinate frame. This allows the computation of the relation between static objects offline and greatly reduces computation. Pioneering Equivariant Planning (PEP) [131] further reduces computation by using center-relative instead of pairwise relative coordinates, i.e. shifting the coordinate frame into the mean position of all processed agents before each transformation. This makes it easier to represent and model interactions. It also eases the integration of prediction and planning, as both can use a shared scene representation. Thus, they can be combined in a single compute-efficient step.
The downside of the Cartesian frames mentioned above is that the models are vulnerable to distributional shifts and often produce inadmissible offroad trajectories [134, 135]. Some works counter this with loss functions that penalize offroad driving [136, 137, 138, 139], driving direction compliance [140] or being distant from the centerline [141].
An alternative approach is to use Frenet representations that decompose positions into progress along and lateral offset from a lane [142, 143, 89, 144]. While this improves map compliance (i.e., reduces offroad), it impairs interaction modeling due to the non-euclidean coordinate frame. Moreover, this coordinate frame cannot be shared across multiple agents.
III-A4 Output Representations
As suggested in Fig. 2, the interfaces between modules are crucial for the overall system performance. To optimize an automated driving system for the final planning task, differentiable interfaces are required [37, 145]. To provide interpretability, the interfaces should comprise a human-readable representation [62]. When predicting trajectories in modular E2E systems, the output must fulfill these two criteria. Additional difficulty arises from the inherent uncertainty of future prediction [146, 147]. Since the intentions of SVs are unknown, multiple future modalities must be considered to make a safe plan [5]. Therefore, the multimodality of $P_{\textrm{pred}}$ is commonly expressed as an explicit density function or estimated with a discrete trajectory set.
@@ III-A4 Output Representations
Explicit density functions model the probability density function of $P_{\textrm{pred}}$ [62, 67, 70, 71]. Continuous distributions like bi-variate Gaussian distributions [148], Gaussian Mixture Models [93, 149, 35, 96], or discrete rasterized heatmaps [150, 89] are used in many prediction modules. Alternatively, object-agnostic representations like occupancy maps [151, 152], or flow fields [153, 154, 71] can be used. Object-agnostic representations handle multimodality implicitly [62] and were shown to be more robust against perturbations [63]. Further, density functions are interpretable for humans. However, since no discrete trajectories are decoded, it is difficult to compare such representations with expert logs for performance evaluation.
@@
In an alternative approach to explicit density functions, discrete trajectory sets are either created by sampling from an intermediate distribution or directly output by the model. To sample a diverse trajectory set from an intermediate distribution, generative models such as Generative Adversarial Networks (GANs) [155], Conditional Variational Autoencoders (CVAEs) [156], normalizing flows [157] or denoising diffusion [95], can be used [8, 158, 159, 160, 161]. Alternatively, recent methods apply several prediction heads to the features extracted by a common backbone to decode a set of diverse trajectories [39, 42, 162, 92]. Other ways to obtain diverse predictions are training loss functions [76], entropy maximization [40, 163], variance-based non-maximum suppression [164], greedy goal sampling [165, 94], divide and conquer strategy [166], evenly spaced goal states [167], ensembling [168], or using pre-defined anchor trajectories [104, 169, 170]. Nevertheless, a guarantee of coverage can only be given by specifying high-level behaviors [171] or regions of the map that must be covered by at least one prediction [77].
III-B Planning
III-B1 Input Representations
The input of the planner depends on the order of modules in sequential IPPSs. If prediction is conditioned on potential EV actions, planning is executed first and takes the sparse or rasterized scene representation (cf. Sec. III-A1) as input. If prediction is executed before planning, the predictions are fed to the planner as an additional input in one of the formats described in Sec. III-A4.
In both cases, goal information for the EV is available in the planning module. It can have different forms like lane-level route information [172], high-level commands [173, 174], and sparse goal points [173]. There are four categories how goal information is incorporated into the IPPS: input features, separate submodules, routing cost, and route attention. Simply providing goal information as input feature yields no guarantees for goal compliance but is straightforward and widely adopted [175, 176, 26, 171, 177, 178, 23, 64, 65, 179, 22, 180]. Separate submodules are used only with high-level commands, which serve as a switch between command-specific submodules [175, 21, 181, 68, 171]. This can prevent the model from over-adapting to driving modes that dominate the training dataset, but requires a fixed number of high-level commands to be defined in advance. A third option to incorporate goal information is by optimizing a cost term, which comprises progress and route compliance [67, 70, 63]. Balancing multiple targets in a cost term provides flexibility and interpretability but can lead to undesirable tradeoffs with safety. Lastly, route attention applies spatial attention to on-route parts of the map. This is achieved by Transformer cross-attention [182] or by removing the off-route portion of the map after initial feature aggregation [183, 131, 184, 185, 38].
@@ III-B1 Input Representations
Since sequential IPPSs consist of discernible modules, the interfaces between them have to comply with diverse requirements. On one hand they should be interpretable to understand failures and facilitate development. Rasterized [56, 68, 171, 186] and sparse [187, 188, 184, 189, 23] scene representations (cf. Sec. III-A1) are teherefore common intermediate representations, which are input to the planning module. Similar to the input representation used by prediction modules, a clear trend can be observed from rasters to sparse inputs [56, 62, 35, 38]. On the other hand an interface should convey all relevant information to downstream modules to cover the long-tailed distribution of potential driving scenarios. Irrespective of the order of modules, information loss is a problem of hand-crafted interfaces [145]. For instance, if a vehicle is currently occluded, but its presence can be inferred from the reflection of its headlight in another car, bounding box representations [56], occupancy fields [71], or affordances [190, 181] cannot propagate this information to downstream modules [145, 37]. Conversely, latent feature representations allow uncertainty propagation, enabling downstream modules to compensate for errors at earlier stages, such as undetected vehicles [62].
III-B2 Planning Paradigms
@@
When taking a closer look at the planning module, three general planning paradigms can be observed. We introduce each paradigm and discuss its suitability for the integration of prediction and planning. Recall that in Eqns. 5 and 6 we defined planning to be a function $f$ mapping from observational inputs $X_{\textrm{EV}}$, $\overline{\textbf{X}}_{\textrm{SV}}$, $I$ and possibly predictions $\overline{\textbf{Y}}_{\textrm{SV}}$ to a trajectory $Y_{\textrm{EV}}$. In the following, we focus on the planning function $f$. To this end, we decompose $f$ into two parts: a proposal generator $g$, yielding a set of multiple potentially suitable trajectories $\hat{Y}^{(i)}_{\textrm{EV}},i=1,..,N_{\textrm{proposals}}$ and a proposal selector $h$ that selects the final plan $Y_{\textrm{EV}}$ among the proposed set. Hence, in this work, we propose to represent the planning function $f$ as
$f=h(g(X_{\textrm{EV}},\overline{\textbf{X}}_{\textrm{SV}},I)$ (7)
or in case of prediction-conditioned planning as
@@
Cost function optimization. As planning aims to find a trajectory that optimizes objectives such as safety, comfort, and progress, it is a natural approach to design and optimize a cost function that balances these potentially conflicting objectives. Cost function-based planning methods [191, 62, 192] rely entirely on the selection function $h$ to find a suitable trajectory,
@@
where $c$ is a cost function of the planned trajectory. Consequently, the proposal generator $g$ may randomly sample feasible motion profiles [62] or obtain them by clustering real-world expert demonstrations [71]. Traditional non-learning-based methods rely on hand-crafted cost functions [193, 194, 195, 196, 197, 198, 199, 200]. However, designing a cost function that effectively balances these goals and generalizes to the long-tailed distribution of possible driving scenarios is challenging. Learning-based methods aim to address this by learning a cost function directly from expert demonstrations [201, 202, 203]. The learned cost function can be non-parametric [62, 204, 205] or comprise hand-designed sub-costs of which only the weights are learned [70, 71, 67, 191]. Cost function-based planning enables the intuitive integration with a prediction module by including predictions in the cost function [69, 70]. Depending on the prediction output, this can be realized either by explicit collision checking [189] or by adopting predicted density functions [62] as non-parametric cost. For example, the NMP [62] predicts a cost volume, which assigns a cost to each potential future EV position within the planning horizon. In contrast, P3 (Perceive, Predict, and Plan) [67] and MP3 (Map, Perceive, Predict and Plan) [71] apply hand-designed sub-costs with learned weighting to heuristically evaluate safety, comfort, traffic rule compliance, and progress of candidate trajectories [191]. In both cases, the cost function is then used to select the min-cost plan from the candidates.
Regression approaches rely entirely on the proposal generator $g$. The selection function $h$ is the identity, and all burden is placed on the proposal generator $g$, which only generates a single proposal $Y_{\textrm{EV}}=\hat{Y}^{(1)}_{\textrm{EV}}$. Typically, regression-based planning is not used in sequential IPPSs but rather found in undirected IPPSs [9, 25, 206, 181, 179, 21, 63, 10, 64, 65, 22, 68, 180, 26, 23, 171, 187]. Unlike cost function-based planning, regression cannot include predictions explicitly. One possibility is to condition the plan on predictions as stated in Eqn. 8. Regression approaches directly output future waypoints  [184, 26, 23] or, more rarely, actions [9, 25, 190, 206]. Both representations can be converted into one another using a kinematic model [49, 55]. Commonly, they are trained by supervised learning from expert data [207], called behavior cloning [208, 209, 210]. This is easy to implement but vulnerable to distributional shifts, which occur when the model reaches states not covered by the training data [23]. Mitigation strategies include heuristically generated data augmentations or policy rollouts during training [23, 211].